OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10? 605 660 748 816

OpenStudy (campbell_st):

you need to find the 1st term \[T _{n} = a+(n-1)d\] using the information for the 12th term \[13 = a + (12 - 1)\times(-10)\] \[13 = a - 110\] I'll let you find a then for the sum use \[s _{12} = \frac{12}{2}[a + 13]\] you found a above..

OpenStudy (anonymous):

a = 123?

OpenStudy (anonymous):

s12 = 738?